40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 287 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  3100   Wed Jun 23 11:25:14 2010 Katharine and SharmilaUpdateWIKI-40M UpdateMaglev

Weekly update


Lab work

We compared the magnetic field strength for 4 magnets in the original setup. The standard deviation was 3.15 G which corresponds to a variation of 2.4%. We had encountered difficulties with the stability of the Gaussmeter. The tip of the Gaussmeter was unsteady and wobbling which led to huge variations for a small change in distance. We stabilized the meter by taping it to a pencil and securing it with wire ties to an aluminum block. We then used translation stages to find the point of maximum field strength for each magnet, which allowed us much more stable readings.

Readings

We are reading and learning about feedback control systems. 

Modelling

Learning to model in Comsol. Our goals for the 1X1 model include incorporating the gravitational force in the measurements and find the distance for which attraction is the strongest, and experimenting with the mesh density and boundary conditions of the domain.

Meetings/seminars

Attended many meetings, including:
Laser safety training
SURF safety training
LIGO seminars
Journal club
LIGO experimental group meeting

  3099   Tue Jun 22 20:07:08 2010 JenneUpdate40m UpgradingFirst attempt at Tip Tilt hanging

[Jenne, Steve, Nancy, Gopal]

We made an attempt at hanging some of the Tip Tilt eddy current dampers today. 

Photo 1 shows the 2 ECDs suspended.

Procedure:

(1) Loosen the #4-40 screws on the side of the ECDs, so the wire can be threaded through the clamps.

(2) Place the ECDs in the locator jigs (not shown), and the locator jigs in the backplane (removed from main TT structure), all laying flat on the table.

(3) Get a length of Tungsten wire (0.007 inch OD = 180um OD), wipe it with acetone, and cut it into 4 ~8cm long segments (long enough to go from the top of the backplane to the bottom).

(4) Thread a length of wire through the clamps on the ECDs, one length going through both ECDs' clamps.

(5) One person hold the wire taught, and straight, and as horizontal as possible, the other person tightened the clamping screws on the ECDs.

(6) Again holding the wire in place, one person put the clamps onto the backplane (the horizontal 'sticks' with 3 screws in them).

(7) The end. In the future, we'll also clip off extra pieces of wire.

When we held up the backplane to check out our handy work, it was clear that the bottom ECD was a much softer pendulum than the top one, since the top one has the wire held above and below, while the bottom one only has the wire held on the top.  I assume we'll trim the wire so that the upper ECD is only held on the top as well?

Lessons learned:

* This may be a 3 person job, or a 2 people who are good at multitasking job.  The wire needs to be held, the ECDs need to be held in place so they don't move during the screwing/clamping process, and the screws need to be tightened.

* Make sure to actually hold the wire taught. This didn't end up happening successfully for the leftmost wire in the photo, and the wire is a bit loose between the 2 ECDs.  This will need to be redone.

* We aren't sure that we have the correct screws for the clamps holding the wire to the backplane.  We only have 3/16" screws, and we aren't getting very many threads into the aluminum of the backplane.  Rana is ordering some 316 Stainless Steel (low magnetism) 1/4" #4-40 screws.  We're going for Stainless because Brass (the screws in the photo), while they passed their RGA scan, aren't really good for the vacuum.  And titanium is very expensive.  

The 2nd photo is of the magnet sticking out of the optic holder.  The hole that the magnet is sitting in has an aluminum piece ~2/3 of the way through.  A steel disk has been placed on one side, and the magnet on the other.  By doing this, we don't need to do any press-fitting (which was a concern whether or not the magnets could withstand that procedure), and we don't need to do any epoxying.  We'll have to wait until the ECDs are hung, and the optic holder suspended, to see whether or not the magnet is sticking out far enough to get to the ECDs. 

Attachment 1: 2_ECDs_small.jpg
2_ECDs_small.jpg
Attachment 2: MagnetStickingOutFar_small.jpg
MagnetStickingOutFar_small.jpg
  3098   Tue Jun 22 18:56:32 2010 JenneUpdateEnvironmentBad placement of recycling bin

Someone has been moving the big blue recycling bin in front of the laser-chiller-chiller (the air conditioner in the control room).  This is unacceptable.  The chiller temp was up to 20.76C.  No good. 

You are free to move the recycling bin around so you can access drawers or the bike-exit-door in the control room, but make sure that it does not block air flow between the chiller-chiller and the chiller. 

The attached photo shows the BAD configuration.

Attachment 1: BlockingLaserChillerChiller_small.jpg
BlockingLaserChillerChiller_small.jpg
  3097   Tue Jun 22 13:38:13 2010 KojiUpdatePhase CameraCurrent phase camera setup. Seeing Acoustic beat

1. In terms of the AOM:

How much beam power is incident on the AOM? How much is the deflection efficiency?
i.e. How much is the power lost by the crystal, deflected in the 1st order, and remaining in the oth order?

I am curious because I assume the AOM (which vender?) is designed for 1064nm and the setup uses 632nm.

2. In terms of the EOM:

How much sidebands do you expect to have?

I assume the EOM is designed for 1064nm, the only difference is the coating at the end. Is this right? 

3. Beating

How much beating strength do you expect?

Is your beating level as expected?

How much is the contrast between the PM sideband and the frequency shifted carrier?
This must include the consideration on the presence of the carrier and the other sidebands.

Quote:

 We've set up a preliminary test bed for the phase camera. It simply uses a HeNe that is split into two beams. One is frequency shifted by an AOM by -40 MHz - df, where df is some acoustic frequency. The second beam is transmitted through a 40MHz EOM to get sidebands. The two beams are recombined and are, currently, incident on a photodetector, but this can be replaced by the phasecamera.

We turned everything on with df = 1kHz and confirmed that a 1kHz signal is visible on the output from the photodetector (PD). The signal looks to be about 1:300 of the DC level from the PD.

 

  3096   Tue Jun 22 09:45:21 2010 Aidan, Joe, RazibUpdatePhase CameraCurrent phase camera setup. Seeing Acoustic beat

 We've set up a preliminary test bed for the phase camera. It simply uses a HeNe that is split into two beams. One is frequency shifted by an AOM by -40 MHz - df, where df is some acoustic frequency. The second beam is transmitted through a 40MHz EOM to get sidebands. The two beams are recombined and are, currently, incident on a photodetector, but this can be replaced by the phasecamera.

We turned everything on with df = 1kHz and confirmed that a 1kHz signal is visible on the output from the photodetector (PD). The signal looks to be about 1:300 of the DC level from the PD.

Attachment 1: 2010-06-22_Phase_camera_layout_version_1.pdf
2010-06-22_Phase_camera_layout_version_1.pdf
  3095   Mon Jun 21 20:11:21 2010 KojiUpdatePhotosInspection of Magnets for the TTs

Were these magnets chipped before the Ni plating?

RA: Yes, it looks like this is the case. We also smashed some of the magnets against a metal surface and saw that a black grime was left. We should hold the magnets with a clean teflon clamp to measure the Gauss. Then we have to wipe the magnets before installing. I share Jenne's concern about the press-fit damaging the plating and so we need to consider using using glue or the ole magnetic attachment method. We should not rely on the structural integrity of the magnets at all.

  3094   Mon Jun 21 18:01:34 2010 JenneUpdate40m UpgradingSRM, PRM hung, magnets inspected

[Jenne, Kiwamu, Rana, Eric Gustafson]

The SRM and PRM have been re-hung, and are ready for installation into the chambers.  Once we put the OSEMs in, we may have to check the rotation about the Z-axis.  That was not confirmed today (which we could do with the microscope on micrometer, or by checking the centering of the magnets in the OSEMs).

Also, Eric and Rana inspected the Tip Tilt magnets, and took a few that they did their best to destroy, and they weren't able to chip the magnets.  There was concern that several of the magnets showed up with the coatings chipped all over the place.  However, since Rana and Eric did their worst, and didn't put any new chips in, we'll just use the ones that don't have chips in them.  Rana confiscated all the ones with obvious bad chips, so we'll check the strengths of the other magnets using a gaussmeter, and choose sets of 4 that are well matched. 

Eric, photographer extraordinaire, will send along the pictures he took, and we'll post them to Picasa.

  3093   Mon Jun 21 14:21:34 2010 Jenne, KiwamuUpdatePhotosInspection of Magnets for the TTs

Some pictures of "magnet inspection" from Picasa.

The coating of some magnets are chipped...

  3092   Sun Jun 20 18:28:25 2010 kiwamuUpdateGreen LockingRe: lock with PDH box

On the wiki I summarized about the modification of the PDH box which is currently running on the end PDH locking. 

http://lhocds.ligo-wa.caltech.edu:8000/40m/Electronics/PDH_Universal_Box

The box was newly labeled "G1" standing for "Green locking #1".

Quote:

by using a modified PDH box the green laser on the X-end station is locked to the arm cavity.

  3091   Sun Jun 20 16:07:23 2010 KojiSummaryCOCCalibration of the metrology lab interferometer

Kiwamu and Koji

Summary

We have visited GariLynn's lab to make a calibration of the metrology interferometer. 

The newly calibrated value is

RoC(SRMU01) = 153.3+/- 1.6 [m]

This is to be compared with the specification of 142m +/- 5m

Although the calibration deviation from the previous value was found to be 1.3%, it is far from explaining the curvature difference between the spec (142m) and the measured value.


Motivation

The previous measurements of the SRM curvatures showed larger RoCs by ~10% compared with the spec.

It can be caused by the mis-calibration of the pixel size of the CCD in the metrology interferometer.
In order to confirm the calibration value, an object with known dimension should be measured by the instrument.

Method

We've got a flat blank optic from "Advanced Thin Film" together with a metalic ring.
The ring has been attached on the blank optic with 3 fragments of a double sided tape.
The RoC of SRMU1 was also measured in order to obtain "the radius of curvature of the day".

The calibration process is as follows:

  1. Measure the diameters of the ring by a caliper in advance to its attachment to the blank.
  2. Determine the inner and outer diameter of the ring in the obtained image.
    Note that the obtained image is pre-calibrated by the default value given by the measurement program
      (i.e. 0.270194mm/pixel for horizontal)
  3. Check the ratio of the diameters with the measured value by the caliper. Correct a systematic effect.
  4. Compare the image measurement and the caliper measurement.

Results

  1. The outer and inner diameters of 2.000" and 1.660" (measured by a caliper, error 0.005"). The ratio is 0.830+/-0.003.
  2. The center and radius for the inner circle were estimated to be (79.7924, 91.6372) and 21.4025 [mm].
    The center and radius for the outer circle were estimated to be (79.6532, 91.6816) and 25.6925 [mm].

    The error would be ~0.01mm considering they sweep 500 pixels by the circle and the pixel size is 0.27mm. i.e. 0.27/Sqrt(500) ~ 0.01mm
  3. Ratio of the inner and outer diameter is 0.8330 +/- 0.0005.
    The systematic error of x is given by solving (21.4025+x)/(25.6925-x)=0.83 ==> x = -0.042 +/- 0.043 [mm]. This is just a 0.2% correction.
    By correcting the above effect, we get (Rin, Rout) = (21.36 +/- 0.046, 25.74 +/- 0.047).
  4. By comparing the result with the caliper measurement, we get calibration factor of 1.013 +/- 0.005.
    This means we measured "1mm" as "1.013mm". The scale was too small.

    We have got the calibration of 0.2737+/-0.0014 [mm/pixel].

Discussion

Because of the calibration error, we measured too long RoC. The same day, we measured the curvature of SRMU01 as 155.26 m.
The newly calibrated value is

RoC(SRMU01) = 153.3+/- 1.6 [m]

This is the value to be compared with the specification of 142m +/- 5m

 

Attachment 1: ring1_inner_centering.pdf
ring1_inner_centering.pdf
Attachment 2: ring1_outer_centering.pdf
ring1_outer_centering.pdf
Attachment 3: SRMU01_pic.png
SRMU01_pic.png
  3090   Sat Jun 19 17:31:48 2010 ranaDAQCDSExcess Noise in C1:IOO-MC_DRUM1 fixed by reboot

I was getting an excess noise in the C1:IOO-MC_DRUM1 channel - it was a flat spectrum of 10 cts/rHz (corresponding to 600 uV/rHz).

I tried a few things, but eventually had to power cycle the crate with c1iovme in order to recover the standard ADC noise level of 3x10^-3 cts/rHz with a 1/sqrt(f) knee at 10 Hz.

I checked the gain of the channel by injecting a 2 Vpp sine wave at 137.035 Hz. 2Vpp as measured on a scope gives 31919 cts instead of the expected 32768, giving a 2.5% error from what we would have naively calculated.

Even so, the noise in this channel is very surprisingly good: 0.003 / (31919 / 2) = 187 nV /rHz.  The best noise I have previously seen from an ICS-110B channel is 800 nV/rHz. What's going on here?

  3089   Fri Jun 18 22:12:29 2010 AlbertoUpdateIOOMode Profile after Mode Matching Telescope (Round 2)

GJ

  3088   Fri Jun 18 21:45:39 2010 kiwamuUpdateIOOMode Profile after Mode Matching Telescope (Round 2)

           [Joe, Kiwamu]

The better mode overlap of 99.3% was achieved by moving MMT2 by ~5 cm 

In the past measurement (elog entry #3077) we already succeeded in getting 99.0% mode overlap.

But according to the calculation there still was a room to improve it by moving MMT2 by 10 cm.

Today we moved MMT2 by ~5 cm which is a reasonable amount we could move because of the narrow space in the chamber.

Eventually it successfully got the better mode overlap.

So we eventually finished mode matching of the new IOOs 


(details)

     Actually moving of MMT2 was done by flipping the mount without moving the pedestal post as Koji suggested. 

At the same time we also flipped the mirror itself (MMT2) so that the curved surface is correctly facing toward the incident beam.

By this trick, we could move the position of MMT2 without losing precious available space for the other optics in the OMC chamber.

     The attached plot shows the result of the mode measurement after the MMT.

During the fitting I neglected the data at x=27 m and 37 m because the beam at those points were almost clipped by the aperture of the beam scan.

- - Here are the fitting results

w0_v             = 2.81183       +/- 7.793e-03  mm  (0.2772%)

w0_h             = 2.9089        +/- 1.998e-02  mm  (0.687%)

z_v             = 5.35487        +/- 0.2244     m   (4.19%)

z_h             = 1.95931        +/- 0.4151     m   (21.18%)

All the distances are calibrated from the position of MMT2 i.e. the position of MMT2 is set to be zero.

        In order to confirm our results, by using the parameters listed above I performed the same calculation of mode overlaps as that posted on the last entry (see here)

The result is shown in Attachement 2. There is an optimum point at x=62mm.

This value is consistent with what we did because we moved MMT2 by ~5 cm instead of 10 cm. 

 

Attachment 1: MMT20100618_edit.png
MMT20100618_edit.png
Attachment 2: newIOO_overlap_edit.png
newIOO_overlap_edit.png
  3087   Fri Jun 18 15:07:26 2010 AlbertoUpdateLSCShort Cavity Length Adjustments

Quote:

You should have been in my lecture yesterday!
Power in the cavity is not a good index (=error signal) to judge the optimal length.
You should look at the phases of the length signals. (i.e. demodulation phase which gives you the maximum amplitude for CARM, PRC, SRC, etc)

You must move the SRC and PRC lengths at the same time.
The resonance of f1 (mostly) depends on the PRC length, but that of f2 depends on both the PRC and SRC lengths. 

Right. Ultimately the phase gain inside the cavity is what we look at. Calculating that for the SBs inside PRC and SRC is actually the first thing I did.

But I kept getting very small angles. Too small, I thought. Maybe there was some problem in the way I calculated it.

Then I made a power analysis to check if the SBs were getting affected at all by that 0.7degree phase shift they're picking up in the arms.

I wanted to show the point where I am, before leaving. But, I keep working on it.

  3086   Fri Jun 18 13:47:20 2010 KojiUpdateLSCShort Cavity Length Adjustments

You should have been in my lecture yesterday!
Power in the cavity is not a good index (=error signal) to judge the optimal length.
You should look at the phases of the length signals. (i.e. demodulation phase which gives you the maximum amplitude for CARM, PRC, SRC, etc)

You must move the SRC and PRC lengths at the same time.
The resonance of f1 (mostly) depends on the PRC length, but that of f2 depends on both the PRC and SRC lengths.

  3085   Fri Jun 18 13:42:52 2010 KojiHowToGeneralUpdate your work

All SURFs (and all others as always) are supposed to post the update of your status on the elog.

In fact, I already heard that Sharmila had been working on the serial connection to TC-200 and made some results. All of us like to hear the story.

  3084   Thu Jun 17 17:09:44 2010 AlbertoUpdateLSCShort Cavity Length Adjustments

I calculated the phase shifts that the sidebands would pick up in the arms in the case we changed the arm length to 38.4m as proposed. I obtained the following values (in degrees):

phi(-f2) = 0.66; phi(-f1) = -0.71; phi(f1) = 0.71; phi(+f2) = -0.66

These are the plots with the results as I obtained from an Optickle simulation (the second zooms in around 38.4m).

sidebandPhaseRotation_73430639654.png sidebandPhaseRotation_73430656054.png

These values agree with what Koji had already estimated (see elog entry 3023).

Since we can't make the arm longer than that, to increase the distance from the resonance, we would like to adjust the length of the short cavities to compensate for that.  For f2 (=55MHz), 0.7 degrees correspond to about 5cm. That is about the length change that we expect to make to the design.

I simulated with Optickle the effect of changing the length of either the SRC or the PRC. The best way I found to do that, was to measure the cavity circulating power when the macroscopic lengths change.

The following plots show the effect of changing either the PRC or SRC length (left or right figure), on the circulating power of both cavities at the same time (top and bottom plots).

shortCavityCirculatingPower_73430666992.png prcCirculatingPower_73430665955.png

 You can compare these with the case of perfect antiresonance as in the following plots:

shortCavityCirculatingPower_73430668892.png shortCavityCirculatingPower_73430669604.png

It seems that the design length for the short cavities are not too bad. f1 is not optimized in the PRC, but changing the length of the cavity wold just make f2 worse in SRC.

These simulations seem to support the choice of not changing the design cavity lengths for PRC and SRC.

Of course these are only an "open loop" simulations. At the moment we don't know what would be the effect of closing the control loops. That is something I'm going to do later. It'll be part of my studies on the effects of cavity absolute length on the whole IFO.

  3083   Wed Jun 16 18:44:07 2010 AlbertoConfigurationComputers40MARS

Quote:

i added my laptop's mac address to teh martian at port 13 today.

 

No personal laptop is allowed to the martian network. Only access to the General Computing Side is permitted.

Please disconnect it.

  3082   Wed Jun 16 18:14:13 2010 AidanUpdateGeneralGlass cover from overhead light smashed on PSL table

I was giving a tour of the 40m yesterday. We were looking at the PSL table. About 30 seconds after I turned the lights on a glass cover from one of the lights (NW corner) popped out of its holder and smashed on the table.

I've cleaned up all the broken glass I could see but there may be some small shards there. Please use caution in that area.

Attachment 1: DSC_1769.JPG
DSC_1769.JPG
Attachment 2: DSC_1768.JPG
DSC_1768.JPG
  3081   Wed Jun 16 18:12:16 2010 nancyConfigurationComputers40MARS

i added my laptop's mac address to teh martian at port 13 today.

 

  3080   Wed Jun 16 11:31:19 2010 josephbSummaryComputersRemoved scaling fonts from medm on Allegra

Because it was driving me crazy while working on the new medm screens for the simulated plant, I went and removed the aliased font entries in /usr/share/X11/fonts/misc/fonts.alias that are associated with medm.  Specifically I removed the lines  starting with widgetDM_.  I made a backup in the same directory called fonts.alias.bak with the old lines.

Medm now behaves the same on op440m, rosalba, and allegra - i.e. it can't find the widgetDM_ scalable fonts and defaults to a legible fixed font.

  3079   Tue Jun 15 21:28:44 2010 albertoUpdateLocking40m Upgrade Optickle Model

Quote:

Quote:

 

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

Carrier and SB (f2) shouldn't be resonant at the same time in the SRC-arms coupled cavity. No additional filtering of the GW signal is wanted.

The SRC macroscopic length is chosen to be = c / f2 - rather than = [ (n+1/2) c / (2*f2) ] - accordingly to that purpose.

I calculated the frequency of the double cavity pole for the 40m SRC-arm coupled cavity.

w_cc = (1 + r_srm)/(1- r_srm) * w_c

where w_c is the arm cavity pole angular frequency [w_c = w_fsr * (1-r_itm * r_etm)/sqrt(r_itm*r_etm) ]

I found the pole at about 160KHz. This number coincides with what I got earlier with my optickle model configured and tuned as I said in my previous entry. See attachments for plots of transfer functions with 0 and 10pm DARM offsets, respectively.

I think  the resonance at about 20 Hz that you can see in the case with non-zero DARM offset, is due to radiation pressure. Koji suggested that I could check the hypothesis by changing either the mirrors' masses or the input power to the interferometer. When I did it frequency and qualty factor of the resonance changed, as you would expect for a radiation pressure effect.

Independently, Jan also calculated the pole frequency of the transfer function DARM / ASQ2 as we would expect it for the SRC-coupled cavity. He also found the pole at about 160KHz. I'm attaching the plot with the transfer function he calculated.
He also said that the little bump at the pole frequency is OK considering that our signal recycling cavity is not much shorter than the arms.

This gave me more confidence about my optickle model of the 40m. This is quite comforting since I used that model other times in the past to calculate several things (i.e. effects of higher unwanted harmonics from the oscillator, or, recently, the power at the ports due to the SB resonating in the arms).

I don't know anymore what Valera said that wasn't right.
Also, as he said, he set it for the carrier to be resonant in the SRC-arms couple cavity. But that is not our case.
Attachment 1: allTransferFunctions_DARMoff_0.pdf
allTransferFunctions_DARMoff_0.pdf allTransferFunctions_DARMoff_0.pdf allTransferFunctions_DARMoff_0.pdf
Attachment 2: allTransferFunctions_DARM2AS_10pmDARMoffset.pdf
allTransferFunctions_DARM2AS_10pmDARMoffset.pdf allTransferFunctions_DARM2AS_10pmDARMoffset.pdf allTransferFunctions_DARM2AS_10pmDARMoffset.pdf
Attachment 3: Jan_DARM2AS.pdf
Jan_DARM2AS.pdf
  3078   Tue Jun 15 19:24:25 2010 ranaUpdateIOOMode Profile after Mode Matching Telescope

Quote:

We obtained a good mode match overlap of 99.0% for the new IOO.

And if we move the position of MMT2 by another 10 cm away from MMT1, we will have 99.6% overlap.

 That's hot stuff.

  3077   Tue Jun 15 16:28:32 2010 kiwamuUpdateIOOMode Profile after Mode Matching Telescope

We obtained a good mode match overlap of 99.0% for the new IOO.

And if we move the position of MMT2 by another 10 cm away from MMT1, we will have 99.6% overlap. 

Yesterday Jenne and I put MMT2 on the OMC table. MMT2 was carefully put by measuring the distance between MMT1 and MMT2.

The position looked almost the same as that drawn on the CAD design.

After putting it we measured the profile after the MMT.

 

The attached figure shows the computed mode overlap according to the fitting result while changing the position of MMT2 in a program code.

The x-axis is the position of MMT2, the current position is set to be zero. The y-axis is the mode match overlap.

Right now the overlap is 99.0% successfully, but this is not an optimum point because the maximum overlap can be achieved at x=100 mm in the plot.

It means we can have 99.6% by moving the position of MMT2 by another 10 cm. This corresponds to an expansion of the MMT length.

If this expansion is difficult due to the narrow available space in the chamber, maybe staying of MMT2 at the current position is fine.

Attachment 1: newIOO_overlap_edit.png
newIOO_overlap_edit.png
  3076   Mon Jun 14 22:16:08 2010 JenneUpdateIOOMode scan after Mode Matching Telescope

[Jenne, Kiwamu]

We measured the mode after the Mode Matching Telescope. 

 

---- fitted parameters ----

       w0_h =  2.85 +/- 0.0115 mm

       w0_v =  2.835 +/- 0.00600 mm

       z0_h = 5.4 +/- 0.447  m

       z0_v  = 6.9 +/- 0.305  m

Attachment 1: after_MMT.png
after_MMT.png
  3075   Mon Jun 14 07:57:07 2010 albertoUpdateLocking40m Upgrade Optickle Model

Quote:

 

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

Carrier and SB (f2) shouldn't be resonant at the same time in the SRC-arms coupled cavity. No additional filtering of the GW signal is wanted.

The SRC macroscopic length is chosen to be = c / f2 - rather than = [ (n+1/2) c / (2*f2) ] - accordingly to that purpose.

  3074   Sun Jun 13 08:28:44 2010 valeraUpdateLocking40m Upgrade Optickle Model

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

  3073   Sat Jun 12 19:43:19 2010 AlbertoUpdateWIKI-40M UpdateIFO modeling Wiki Page updated

Today I started writing the IFO modeling wiki page.

The idea is to make it a reference place where to share our modeling tools for the 40m.

  3072   Sat Jun 12 19:41:04 2010 AlbertoUpdateLocking40m Upgrade Optickle Model

I wrote down the settings according to which I tuned the optickle model of the 40m Upgrade.

Basically I set it so that:

  1. PRC alone anti-resonant for the carrier and resonant for both sidebands
  2. SRC alone resonant for the carrier and resonant for the f2 sideband

In this way when the carrier becomes resonant in the arms we have:

  1. carrier resonant in PRC and anti-resonant in SRC
  2. f1 resonant in PRC and non resonant in SRC
  3. f2 resonant in SRC

The DARM offset for DC readout is optional, and doesn't change those conditions.

I also plotted the carrier and the sideband's circulating power for both recycling cavities.

I'm attaching a file containing more detailed explanations of what I said above. It also contains the plots of field powers, and transfer functions from DARM to the dark port. I think they don't look quite right. There seems to be something wrong.

Valera thought of fixing the problem, removing the 180 degree offset on the SRM, which is what makes the sideband rather than the carrier resonant in SRC. In his model the carrier becomes resonant and the sideband anti-resonant. I don't think that is correct.

The resonant-carrier case is also included in the attachment (the plots with SRMoff=0 deg). In the plots the DARM offset is always zero.

I'm not sure why the settings are not producing the expected transfer functions.

Attachment 1: optickleIFOworkingpoint.pdf
optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf optickleIFOworkingpoint.pdf
  3071   Sat Jun 12 18:03:00 2010 sharmilaUpdateelogTemperature Controller

Kiwamu and I setup a serial port terminal for receiving data from TC200 via a RS-232 USB interface. It was done using a Python code. Some command definitions need to be done to get the output from TC-200.

  3070   Fri Jun 11 22:09:58 2010 valeraHowToCDSfoton

 It appears that foton does not like the unstable poles, which we need to model the transfer functions.

But one can try to load the filters into the front end by generating the filter file e.g.:

#
# MODULES DARM_ASDC
 
#
################################################################################
### DARM_ASDC                                                                   ###
################################################################################
# SAMPLING DARM_ASDC  16384
# DESIGN   DARM_ASDC  
### ####
DARM_ASDC  0 21 6  0  0 darm 1014223594.005454063416 -1.95554205062071  0.94952075557861 0.06176931505784 -0.93823068494216
                         -2.05077577179611   1.05077843532639  -2.05854170261687  1.05854477394411
                         -1.85353637553024   0.86042048250739  -1.99996540107622  0.99996542454814 
                         -1.93464836371852   0.94008893626414  -1.89722830906561  0.90024221050918
                         -2.04422931770060   1.04652211283968  -2.01120153956052  1.01152717233685 
                         -1.99996545575365   0.99996548582538  -1.99996545573320  0.99996548582538

 

 

 

Unfortunately if you open and later save this file with foton it will strip the lhp poles.

  3069   Fri Jun 11 15:04:25 2010 josephbUpdateCDSMulti-filter matrix medm screens finished and script for copying filters from SOS file

I've finished the MEDM portion of the RCG FiltMuxMatrix part.  Now it generates an appropriate medm screen for the matrix, with links to all the filter banks.  The filter bank .adl files are also generated, and placed in a sub directory with the name of the filter matrix as the name of the sub directory.

The input is the first number and the output is the second number.  This particular matrix has 5 inputs (0 through 4) and 15 outputs (0 through 14). Unfortunately, the filter names can't be longer than 24 characters, which forced me to use numbers instead of actual part names for the input and output.

 

The key to the numbers is:

Inputs:

DARM 0
MICH 1
PRC 2
SRC 3
CARM 4

Outputs:

AS_DC 0
AS11_I 1
AS11_Q 2
AS55_I 3
AS55_Q 4
POP_DC 5
POP11_I 6
POP11_Q 7
POP55_I 8
POP55_Q 9
REFL_DC 10
REFL11_I 11
REFL11_Q 12
REFL55_I 13
REFL55_Q 14

To get this working required modifications to the feCodeGen.pl and the creation of mkfiltmatrix.pl (which was based off of mkmatrix.pl).  These are located in /cvs/cds/caltech/cds/advLigoRTS/src/epics/util/

In related news, I asked Valera if he could load the simulated plant filters he had generated, and after several tries, his answer was no.  He says it has the same format as those filter they pass to the feed forward banks down in Livingston, so he's not sure why they won't work.

I tested my script, FillFotonFilterMatrix.py, on some simple second order section filters (like gain of 1, with b1 = 1.01, b2 = 0.02, a1 = 1.03, a2 = 1.04), and it populated the foton filter file correctly, and was parsed fine by Foton itself.  So I'm going to claim the script is done and its the fault of the filters we're trying to load.  This script is now living in /cvs/cds/caltech/chans/ , along with a name file called lsp_dof2pd_mtrx.txt which tells the script that DARM is 0, CARM is 1, etc.  To run it, you also need a SOS.txt file with the filters to load, similar to the one Valera posted here, but preferably loadable.

I also updated my progess on the wiki, here.

Attachment 1: FiltMatrixMedm.png
FiltMatrixMedm.png
  3068   Fri Jun 11 14:31:04 2010 kiwamuUpdateIOOmode matching of new IOO

We decided not to care about the mode after MMT1.

So far Koji, Alberto and I have measured the beam profile after MMT1,

but we are going to stop this measurement and go ahead to the next step i.e. putting MMT2

There are two reasons why we don't care about the profile after MMT1:

     (1) it is difficult to fit the measured data

     (2) The position of MMT1 is not critical for the mode matching to the IFO.

The details are below.

 


(1) difficulty in fitting the data

The precision of each measured point looked good enough, but the fitting result varies every measurement.

The below shows the data and their fitted curves. 

 profile_MMT.png

In the label, "h" and "v" stand for "horizontal" and "vertical" respectively.

The solid curves represent the fitting results, varying by each measurement.

In order to increase the reliability of the fitting, we had to take some more data at further distance.

But we couldn't do it, because the beam radius already becomes 3 mm even at 2 m away from MMT1 and at this point it starts to be clipped on the aperture of the beam scan.

Thus it is difficult to increase the reliability of the fitting. 

Once if we put MMT2 the beam should have a long Rayleigh range, it means we can measure the profile at further distance, and the fitting must be more reliable.

 


(2) positioning of MMTs

Actually the position of MMT1 is not so critical for the mode matching. 

The most important point is the separation distance of MMT1 and MMT2.

As written in Jenne's document, if we slide the positions of MMT1 and MMT2 while keeping their appropriate separation distance, the mode match overlap still stays above 99%

This is because the beam coming from MC3 is almost collimated (ZR~8m), so the position of MMTs doesn't so matter. 

To confirm it for the real case, I also computed the mode overlap while sliding the position of MMTs by using real data. The below is the computed result.

dist_MMO.png

It is computed by using the measured profile after MC3 (see the past entry).

The overlap still stay above 99% when the distance from MC to MMT is between 1300 and 3000mm.

This result suggests to us putting MMT1 as we like.  

  3066   Fri Jun 11 13:32:28 2010 KojiUpdateGreen Lockingend PDH with thermal feedback

GJ!

Quote:

A thermal feedback was installed to the end PDH locking and it works well. There are no saturations 

As I said the feedback signal was sometimes saturated at the sum-amp because the drive signal going to the laser PZT was large at low frequency (below 1Hz).

So I made a passive low pass filter which filters the signal controlling the temperature of the laser crystal, and put it before the temperature drive input.

Now the amount of the feedback signal got reduced when it is locked, and there are no saturations. It's very good.

 

  3065   Fri Jun 11 11:54:42 2010 kiwamuUpdateGreen Lockingend PDH with thermal feedback

A thermal feedback was installed to the end PDH locking and it works well. There are no saturations 

As I said the feedback signal was sometimes saturated at the sum-amp because the drive signal going to the laser PZT was large at low frequency (below 1Hz).

So I made a passive low pass filter which filters the signal controlling the temperature of the laser crystal, and put it before the temperature drive input.

Now the amount of the feedback signal got reduced when it is locked, and there are no saturations. It's very good.


(thermal property of the crystal) 

According to the specification sheet for the 1W Innolight, the thermal properties of the crystal are:

  Response coefficient : 3GHz/K 

  Temperature control coefficient : 1K/V

  Thermal response bandwidth: 1Hz 

 

(filter circuit and actuator response)

In order to feedback the signal blow 1Hz, a low pass fiter is needed. 

The attachment:1 shows the filter circuit I made.

Since I found that the drive input had an input impedance of 100kOhm, so I put relatively big resistors to have a moderate gain.

The expected actuator responses are also attached.

The blue curve represents the response of the PZT, the green is the thermal response including the low pass filter and the red curve is the total response composed of both the responses.

I assume that the PZT response is 1MHz/V according to Mott's measurement.

Also I assume that the thermal response intrinsically has two poles at 1Hz according to the specification listed above.

In the total response, there is a little gain reduction around 2Hz due to the cancelation of each other, but it still looks okay.

 

Attachment 1: LPF.png
LPF.png
Attachment 2: thermal_feedback.png
thermal_feedback.png
  3064   Thu Jun 10 11:10:21 2010 AlbertoUpdateGeneralLaTeXlabs

Quote:

I could not dare to share my google doc with this site...

Quote:

Quote:

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

 so cool!

 

Just in case,  granted access to Google docs can be revoked any time from here:

https://www.google.com/accounts/IssuedAuthSubTokens

  3063   Thu Jun 10 10:58:02 2010 KojiUpdateGeneralLaTeXlabs

I could not dare to share my google doc with this site...

Quote:

Quote:

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

 so cool!

 

  3062   Thu Jun 10 07:53:14 2010 AlbertoUpdatePEMLaTeXlabs

Quote:

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

 so cool!

  3061   Wed Jun 9 21:05:44 2010 ranaSummaryComputersop540m is not to be used

This is a reminder (mainly for Steve, who somehow doesn't believe these things) that op540m is not to be used for your general pleasure.

No web, no dataviewer, no DTT. Using these things often makes the graphical X-Windows crash. I have had to restart the StripTool, our seismic BLRMS and our Alarms many times because someone uses op540m, makes it crash, and then does not restart the processes.

Stop breaking op540m, Steve!

  3060   Wed Jun 9 19:47:08 2010 nancyUpdatePEMlead balls on concrete

Quote:

Quote:

Valera and I put the 2 Guralps and the Ranger onto the big granite slab and then put the new big yellow foam box on top of it.

There is a problem with the setup. I believe that the lead balls under the slab are not sitting right. We need to cut out the tile so the thing sits directly on some steel inserts.

You can see from the dataviewer trend that the horizontal directions got a lot noisier as soon as we put the things on the slab.

 The tiles were cut out in 1.5" ID circle to insure that the 7/16" OD lead balls would not touch the tiles on Wednesday, May 26, 2010

Granite surface plate specifications: grade B, 18" x 24" x 3" , 139 lbs

These balls and granite plate were removed by  Rana in entry log #3018 at 5-31-2010

 I tried to calculate the frequency of resonance using Rayleigh's method.  approximated the geometry of lead to be that of a perfect cylinder, and  the deformation in the lead by the deflection in a cantilever under  a shear strain.

this rough calculation gives an answer of 170Hz and depends on the dimensions of each lead, number of leads, and mass of the granite. But the flaw pointed out is that this calculation doesnot depend on the dimension of the granite slab, nor on the exact placing of the lead spheres with respect toteh COM of the slab.

I will put up the calculations details later, and also try to do a FEM analysis of the problem.

 

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

  3059   Wed Jun 9 11:13:11 2010 kiwamuUpdateGreen Lockinglock with PDH box

A progress on the end PDH locking :

by using a modified PDH box the green laser on the X-end station is locked to the arm cavity.

So far the end PDH locking had been achieved by using SR560s, but it was not sophisticated filter.

To have a sophisticated filter and make the control loop more stable, the PDH box labeled "#G1" was installed instead of the SR560s.

After the installation the loop looks more stable than the before.

Some details about the modification of the PDH box will be posted later.

 

Although, sometimes the loop was unlocked because the sum-amp (still SR560) which mixes the modulation and the feedback signal going to the NPRO PZT was saturated sometimes.

Thus as we expected a temperature control for the laser crystal is definitely needed in order to reduce such big low frequency drive signal to the PZT.

  3058   Wed Jun 9 10:31:13 2010 steveUpdateSAFETYsafety training

Nancy and Sharmila received introductory early bird surf safety training for the 40m lab.

  3057   Tue Jun 8 20:52:25 2010 josephbUpdatePEMDAQ up (for the moment)

As a test, I did a remote reboot of both Megatron and c1iscex, to make sure there was no code running that might interfere with the dataviewer.  Megatron is behind a firewall, so I don't see how it could be interfering with the frame builder.  c1iscex was only running a test module from earlier today when I was testing the multi-filter matrix part.  No daqd or similar processes were running on this machine either, but it is not behind a firewall at the moment. 

Neither of these seemed to affect the lack of past data.  I note the error message from dataviewer was "read(); errno=9".

Going to the frame builder machine, I ran dmesg.  I get some disturbing messages from May 26th and June 7th. There are 6-7 of these pairs of lines for each of these days, spread over the course of about 30 minutes.

Jun 7 14:05:09 fb ufs: [ID 213553 kern.notice] NOTICE: realloccg /: file system full

Jun 7 14:11:14 fb last message repeated 19 times

There's also one:

Jun 7 13:35:14 fb syslogd: /usr/controls/main_daqd.log: No space left on device

I went to /usr/controls/ and looked at the file.  I couldn't read it with less, it errored with Value too large for defined data type.  Turns out the file was 2.3 G.  And had not been updated since June 7th.  There were also a bunch of core dump files from May 25th, and a few more recent.  However the ones from May 25th were somewhat large, half a gig each or so.  I decided to delete the main_daqd.log file as well as the core files.

This seems to have fixed the data history for the moment (at least with one 16k channel I tested quickly). However, I'm now investigating why that log file seems to have filled up, and see if we can prevent this in the future.

Quote:

As before, I am unable to get data from the past. With DTT on Allegra I got data from now, but its unavailable from 1 hour ago. Same problem using mDV on mafalda. I blame Joe again - or the military industrial complex.

 

Quote:

Quote:

 Although trends are available, I am unable to get any full data from in the past (using DTT or DV). I started the FB's daqd process a few times, but no luck. 

I blame Joe's SimPlant monkeying from earlier today for lack of a better candidate. I checked and the frames are actually on the FB disk, so its something else.

 I tried running dataviewer and dtt this morning.  Dataviewer seemed to be working.  I was able to get trends, full data on a 2k channel (seismic channels) and full data on a 16k channel (C1:PEM-AUDIO_MIC1)  This was tried for a period 24 hours a go for a 10 minute stretch.

I also tried dtt and was able to get 2k and 16k channel data, for example C1:PEM-AUDIO_MIC1.  Was this problem fixed by someone last night or did time somehow fix it?

 

  3056   Tue Jun 8 18:39:36 2010 ranaUpdatePEMDAQ down

As before, I am unable to get data from the past. With DTT on Allegra I got data from now, but its unavailable from 1 hour ago. Same problem using mDV on mafalda. I blame Joe again - or the military industrial complex.

 

Quote:

Quote:

 Although trends are available, I am unable to get any full data from in the past (using DTT or DV). I started the FB's daqd process a few times, but no luck. 

I blame Joe's SimPlant monkeying from earlier today for lack of a better candidate. I checked and the frames are actually on the FB disk, so its something else.

 I tried running dataviewer and dtt this morning.  Dataviewer seemed to be working.  I was able to get trends, full data on a 2k channel (seismic channels) and full data on a 16k channel (C1:PEM-AUDIO_MIC1)  This was tried for a period 24 hours a go for a 10 minute stretch.

I also tried dtt and was able to get 2k and 16k channel data, for example C1:PEM-AUDIO_MIC1.  Was this problem fixed by someone last night or did time somehow fix it?

  3055   Tue Jun 8 15:58:25 2010 josephb, alexUpdateCDSNew multi-filter matrix part added to RCG (at the 40m at least)

A new webview of the LSP model is available at:

https://nodus.ligo.caltech.edu:30889/FE/lsp_slwebview_files/

This model include a couple example noise generators as well as the new Matrix of Filter banks (5 inputs x 15 outputs = 75 Filters!).  The attached png shows where these parts can be found in the CDS_PARTS library.  I'm still working on the automatic generation of the matrix and filter bank medm screens for this part.  The plan is to have a matrix screen similar to current ones, except that the value entry points to the gain setting of the associated filter.  In addition, underneath each value, there will be a link to the full filter bank screen.  Ideally, I'd like to have the filter adl files located in a sub-directory of the system, to keep clutter down.

I've cut and past the new Foton file generated by the LSP model below.  The first number following the MTRX is the input the filter is taking data from and the second number is the output its pushing data to.  This means for the script parsing Valera's transfer functions, I need to input which channel corresponds to which number, such as DARM = 0, MICH = 1, etc.  So the next step is to write this script and populate the filter banks in this file.

# FILTERS FOR ONLINE SYSTEM
#
# Computer generated file: DO NOT EDIT
#
# MODULES DOF2PD_AS11I DOF2PD_AS11Q DOF2PD_AS55I DOF2PD_AS55Q
# MODULES DOF2PD_ASDC DOF2PD_POP11I DOF2PD_POP11Q DOF2PD_POP55I
# MODULES DOF2PD_POP55Q DOF2PD_POPDC DOF2PD_REFL11I DOF2PD_REFL11Q
# MODULES DOF2PD_REFL55I DOF2PD_REFL55Q DOF2PD_REFLDC Mirror2DOF_f2x1
# MODULES Mirror2DOF_f2x2 Mirror2DOF_f2x3 Mirror2DOF_f2x4 Mirror2DOF_f2x5
# MODULES Mirror2DOF_f2x6 Mirror2DOF_f2x7 DOF2PD_MTRX_0_0 DOF2PD_MTRX_0_1
# MODULES DOF2PD_MTRX_0_2 DOF2PD_MTRX_0_3 DOF2PD_MTRX_0_4 DOF2PD_MTRX_0_5
# MODULES DOF2PD_MTRX_0_6 DOF2PD_MTRX_0_7 DOF2PD_MTRX_0_8 DOF2PD_MTRX_0_9
# MODULES DOF2PD_MTRX_0_10 DOF2PD_MTRX_0_11 DOF2PD_MTRX_0_12 DOF2PD_MTRX_0_13
# MODULES DOF2PD_MTRX_0_14 DOF2PD_MTRX_1_0 DOF2PD_MTRX_1_1 DOF2PD_MTRX_1_2
# MODULES DOF2PD_MTRX_1_3 DOF2PD_MTRX_1_4 DOF2PD_MTRX_1_5 DOF2PD_MTRX_1_6
# MODULES DOF2PD_MTRX_1_7 DOF2PD_MTRX_1_8 DOF2PD_MTRX_1_9 DOF2PD_MTRX_1_10
# MODULES DOF2PD_MTRX_1_11 DOF2PD_MTRX_1_12 DOF2PD_MTRX_1_13 DOF2PD_MTRX_1_14
# MODULES DOF2PD_MTRX_2_0 DOF2PD_MTRX_2_1 DOF2PD_MTRX_2_2 DOF2PD_MTRX_2_3
# MODULES DOF2PD_MTRX_2_4 DOF2PD_MTRX_2_5 DOF2PD_MTRX_2_6 DOF2PD_MTRX_2_7
# MODULES DOF2PD_MTRX_2_8 DOF2PD_MTRX_2_9 DOF2PD_MTRX_2_10 DOF2PD_MTRX_2_11
# MODULES DOF2PD_MTRX_2_12 DOF2PD_MTRX_2_13 DOF2PD_MTRX_2_14 DOF2PD_MTRX_3_0
# MODULES DOF2PD_MTRX_3_1 DOF2PD_MTRX_3_2 DOF2PD_MTRX_3_3 DOF2PD_MTRX_3_4
# MODULES DOF2PD_MTRX_3_5 DOF2PD_MTRX_3_6 DOF2PD_MTRX_3_7 DOF2PD_MTRX_3_8
# MODULES DOF2PD_MTRX_3_9 DOF2PD_MTRX_3_10 DOF2PD_MTRX_3_11 DOF2PD_MTRX_3_12
# MODULES DOF2PD_MTRX_3_13 DOF2PD_MTRX_3_14 DOF2PD_MTRX_4_0 DOF2PD_MTRX_4_1
# MODULES DOF2PD_MTRX_4_2 DOF2PD_MTRX_4_3 DOF2PD_MTRX_4_4 DOF2PD_MTRX_4_5
# MODULES DOF2PD_MTRX_4_6 DOF2PD_MTRX_4_7 DOF2PD_MTRX_4_8 DOF2PD_MTRX_4_9
# MODULES DOF2PD_MTRX_4_10 DOF2PD_MTRX_4_11 DOF2PD_MTRX_4_12 DOF2PD_MTRX_4_13
# MODULES DOF2PD_MTRX_4_14  
# MODULES 

Attachment 1: CDS_Library.png
CDS_Library.png
  3054   Tue Jun 8 00:38:22 2010 Koji, KiwamuUpdateIOOimproved Gaussian beam in new IOO

The shape of the beam spot in the new input optics got much much better 

As Alberto and Kiwamu found on the last week, the beam spot after MMT1 had not been good. So far we postponed the mode measurement due to this bad beam profile.

Today after we did several things in the vacuum chamber, the beam spot became really a good Gaussian spot. See the attachment below.

There were two problems which had caused the bad profile:

(1)  a steering mirror after MMT1 with the incident angle of non 45 deg

(2) clipping at the Faraday.

 

Also MCT_QPD and MCT_CCD were recovered from misalignment  

Tomorrow we are going to restart the mode matching. 

 


(what we did)

* We started from checking the shape of the beam going out from the BS chamber. There still were some stripes which looked like an interference on the spot. 

* We found a steering mirror after MMT1 had the incident angle of non 45 deg. In fact the mirror had a large transmission. After we made the angle roughly 45 deg, the stripes disappeared.

However the spot still didn't look a good Gaussian, it looked slightly having a bump on the horizontal profile.

* Prior to moving of some optics in the vacuum, we ran the A2L_MC scripts in order to check the beam axis. And it was okay.

* To recover the MCT, we steered one of the vacuum mirrors which was located after the pick off mirror.  And after aligning some optics on the AP table, finally we got MCT recovered.

 * We rearranged MC_refl mirrors according to the new optical layout that Koji has made. At the same time the mirrors for IFO_refl was also rearranged coarsely.

 * We leveled the optical table of the MC chamber by moving some weights. Then we locked the MC again and aligned it. We again confirmed that the beam axis was still fine by running the A2L scripts.

 * We found the beam going through Faraday was off-centered by ~5mm toward the west. So we moved it so that the beam propagates on the center of it. 

 * Then looking at the beam profile after MMT1, we found that the profile became really nicer. It showed a beautiful Gaussian. 

In the attachment below, the top panel represents the horizontal profile and the bottom one represents the vertical profile.

The blue curves overlaid on the plot are fitted Gaussian profile, showing beautiful agreements with the measured profile.

Attachment 1: 2010-6-7_2.png
2010-6-7_2.png
  3053   Mon Jun 7 07:39:38 2010 AlbertoOmnistructureElectronicsCapacitor Bridge Test

Quote:

To get a feel for the Capacitive Bridge problems, we setup a simple bridge using fixed (1 nF) caps on a breadboard. We used an SR830 Lock-In amplifier to drive it and readout the noise.

The measurement setup for the Capacitor Bridge Test is still sitting on one of the work benches.

Unless the experiment is supposed to continue today, the equipment shouldn't have been left on the bench. It should have been  taken back to the lab.

Also the cart with HP network analyzer used for the test was left in the desk area. That shouldn't have left floating around in the desk area anyway.

The people responsible for that, are kindly invited to clean up after themselves.

  3052   Sun Jun 6 08:08:05 2010 rana, sanjitSummaryElectronicsCapacitor Bridge Test

To get a feel for the Capacitive Bridge problems, we setup a simple bridge using fixed (1 nF) caps on a breadboard. We used an SR830 Lock-In amplifier to drive it and readout the noise.

CapacitanceBridge.png

We measured the cap values with an LCR meter. They were all within a few % of 0.99 nF.

With a 0.5 V drive to the top of the bridge, the A-B voltage was ~2 mV as expected from the matching of the capacitors.

(** Note about the gain in the SR830: In order to find the magnitude of the input referred signal, one has to divide by G. G = (10 V)/ Sensitivity. 'Sensitivity' is the setting on the front panel.)

  1. Directly measuring from Vs to ground gives 0.5 V, as expected. This is done to verify the calibration later on.
  2. Shorting the A and B wires to ground gives ~0 V and lets us measure the noise. On the spectrum analyzer it was ~400 nV/rHz at 100 Hz and rising slowly to 4 uV/rHz at 100 mHz. In this state, the sensitivity was 10 mV, so the overall gain was 1000. That gives an input referred level of ~0.4 nV/rHz at the input.
  3. Hooking up now to A-B: the signal is ~10x larger than the 'dark' noise everywhere. 2 uV/rHz @ 100 Hz, 10 uV/rHz @ 10 Hz, 50 uV/rHz @ 1 Hz. The spectrum is very non-stationary; changing by factors of several up and down between averages. Probably a problem with the cheapo contacts in the breadboard + wind. The gain in this state was still 1000. So at 1 Hz, its 50 nV/rHz referred to the input.

To convert into units of capacitance fluctuation, we multiply by the capacitance of the capacitors (1 nF) and divide out by the peak-peak voltage (1 V). So the bridge sensitivity is 50e-9 * 1e-9 = 5 x 10^-17 F/rHz.

If we assume that we will have a capacitive displacement transducer giving 1 nF capacitance change for a 0.1 mm displacement, this bridge would have a sensitivity of 5 x 10^-12 m/rHz @ 1 Hz. We would like to do ~50-100x better than this. The next steps should be:

  1. Solder it all together on a PCB to have less air current sensitivity and decent contacts.
  2. Use a low-noise FET input. Since the impedance of the bridge is ~5 kOhms at this frequency, we are probably current noise limited.
  3. Estimate the oscillator amplitude noise sensitivity.
  3051   Sun Jun 6 04:48:41 2010 ranaUpdateCOCITM01 HR Phase Map

While trying to set up the SIS-FFT to use our new ITM phase maps, I noticed that the surface of our ITMs looks pretty good actually (even compared to the aLIGO pathfinder optic map on the AIC wiki). I'm attaching it here for your viewing pleasure.

The code to plot it has been added to the SVN in the PhaseMaps/mat directory.

Attachment 1: itm01hr.png
itm01hr.png
  3050   Fri Jun 4 23:52:57 2010 ranaUpdatePEMseismometers off of linoleum floor

Untitled.pnghuddlez.png

For the huddle test, I have updated the code to divide the residual by sqrt(2) because of the assumption of equal noise from the 2 Guralps. We would have to multiply this trace by sqrt(2) to compare with the previous results.

Now the question is, how do I add a low noise ~50 mV offset to the front of the Guralp breakout box to test for the noise of the box?

ELOG V3.1.3-